# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

### v2

@@ 5 Empirical validation
In this section we demonstrate the utility of our bounds on both synthetic and real-world data sets. We study three classes of task environments, with very different types of shifts between tasks, and show, that for sufficiently many tasks, the bounds are tight.
Since Corollary 3.4 applies uniformly over all posteriors, it can be optimized at each time step w.r.t. the current posterior. Specifically, we select
$\hat{Q}_{1:i}=\mathop{\mathrm{argmin}}_{Q_{1:i}}\left\{\hat{\mathcal{L}}(Q_{1:i},S_{i})+\frac{1}{\lambda_{i}}D_{\mathrm{KL}}(Q_{1:i}||Q_{1:i-1})\right\}$ (12)
This optimization process yields a practical continual learning algorithm that we can easily apply our upper bound on, detailed in Algorithm 1.
@@
4:  for each task $t$ from $1$ to $T$ do
@@
This algorithm can be implemented in practice via variational inference (VI) using multivariate Gaussian prior and posterior distributions on model parameters. The posterior can be used to estimate the forgetting and its upper bound by using a pre-defined test set as a proxy for the data distribution $\mathcal{D}_{i}$. Calculating the upper bound (3.4) would add an additional calculation of $O(t\cdot(m+d))$ to each task, where $m=|S_{t}|$ and $d$ is the number of model parameters. In total, calculating the bound at each task will add $O(T^{2}\cdot(m+d))$. Compared to standard VI methods, the bound does not require additional memory or samples (it does require sampling from previous test tasks for calculating the bound, but this is also required for calculating forgetting in general).
@@
In order to measure the bound tightness of Algorithm 1, we must define the relevant metrics we wish to measure. These are the backwards transfer $BWT(\hat{Q}_{1:T},[T])$ (measured via a separate test set per task) and the forgetting $F(\hat{Q}_{1:T},[T])$, as well as the average forward transfer $\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}(\hat{Q}_{1:t},\mathcal{D}_{t})$. We compare Algorithm 1 to the expected maximum likelihood estimator (MLE) method, a baseline that only minimizes ^$\hat{\mathcal{L}}(Q_{1:t},S_{t})$ at each task $t$. Naturally, we would also like to consider the upper bound on both the forgetting and backward transfer.
In the following experiments, we use fully connected neural networks with a separate linear head per task to facilitate measuring the forgetting without re-training linear heads. We use Adam (Kingma and Ba, 2015) for optimization. The full list of hyper-parameters is listed in B.
@@
We begin by examining several setups of binary classification tasks in $\mathbb{R}^{10}$. All tasks draw samples from a $10$-dimensional Gaussian distribution $p(x)=\mathcal{N}(x;0,I_{10})$, and $y=\mathrm{sgn}(a^{\top}x)$. To simplify, only the first two features were used to determine $y$, meaning that there is a $2d$ linear separator embedded in $\mathbb{R}^{10}$. We consider the following settings: (1) Similar tasks: linear separators for all tasks are within $10^{\circ}$ of some reference angle. (2) Gradual shift: tasks change angle gradually in a set direction, with each consecutive task being within $10^{\circ}$ of the last one. (3) Orthogonal shift: Tasks come from two distinct, orthogonal angles, with the first half of tasks being from the first angle and the latter half corresponding to the second angle. Two additional settings are considered in Appendix B.
We used a total of $T=100$ tasks in this domain in order to observe general trends. We note that these problems differ significantly in task order and overall behavior, and are aimed to provide a diverse set of challenges. In addition, we should not expect a single algorithm to perform optimally across all settings, as known empirical methods based on regularization would likely perform poorly on orthogonal tasks whereas replay-based and architecture-based methods would have issues with gradual shift. We expect that in real-world settings, prior knowledge can be used to suggest specific algorithms. In particular, the settings of Gradual shift and Orthogonal shifts are constructed such that forgetting may be desirable. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially given changing environments. Of our constructed settings, in both the gradual and orthogonal shift settings, not all previous knowledge is beneficial.
@@
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
@@
MLE $5.4\pm 0.5$ $7.2\pm 0.4$ $-10.9\pm 0.5$ $-9.1\pm 0.8$ $3.5\pm 0.3$
PAC-Bayes $\textbf{3.2}\pm 0.2$ $\textbf{5.3}\pm 0.3$ $-11.5\pm 1.0$ $-9.4\pm 1.1$ $\textbf{2.0}\pm 0.2$
@@
MLE $4.3\pm 0.3$ $6.4\pm 0.4$ $-10.1\pm 0.5$ $-8.0\pm 0.6$ $3.4\pm 0.2$
PAC-Bayes $\textbf{2.8}\pm 0.3$ $\textbf{4.8}\pm 0.3$ $-10.6\pm 0.8$ $-8.6\pm 0.8$ $\textbf{1.9}\pm 0.4$
@@
MLE $11.7\pm 2.7$ $13.9\pm 3.0$ $-6.8\pm 0.8$ $-4.6\pm 1.2$ $7.4\pm 2.5$
PAC-Bayes $\textbf{10.2}\pm 2.4$ $\textbf{12.0}\pm 2.3$ $-7.4\pm 1.3$ $-5.6\pm 1.8$ $\textbf{5.8}\pm 2.7$
@@
(b) Orthogonal shift
Figure 1: Forgetting and upper bound (3.5) over time for $10d$ data. (a) Gradual shift (b) Task shift at $T=50$.
Table 1 describes the forward and backward transfer as well as forgetting for the $10d$ tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign. Best result for each setting is in bold. We can see that for most settings, Algorithm 1 tends to provide good forward transfer, as well as backward transfer and forgetting. For the orthogonal shift setting, we can see that adding a prior for regularization leads to looser upper bounds on both backward transfer and forgetting. Figure 1 compares the forgetting for the gradual shift setting compared to the more abrupt orthogonal shift at $T=50$. We can clearly see that the forgetting upper bound for the PAC-Bayesian method increases abruptly at the point of task shift, as both the KL-divergence and the discrepancy term increase. Due to the sudden shift, the model from the previous task serves as a poor prior and the model is forced to change significantly. We also note that across all settings, the PAC-Bayesian algorithm has lower standard-error and is more stable across different random seeds.
Notably, the final forgetting at $T=100$ as well as its upper bound are negative for all settings. The forgetting itself is negative and seems to follow a trend of increasing at the start, decreasing as general model performance improves and finally increasing towards zero forgetting. This trend is more pronounced in the orthogonal shift setting but occurs to a lesser extent in the other settings. The upper bound (3.4) decreases gradually starting from a positive value to a negative one. We observe that as the shared model improves, forgetting decreases, likely due to the learned shared representation improving overall across several tasks, and we see the forgetting trend towards zero.
An interesting observation is that the backward transfer (and forgetting) for the orthogonal shift setting is much higher. In this setting, forgetting may be a desired phenomenon, as the data distribution shifts in the middle of the learning process. We see that indeed this shift leads to higher forgetting. Since the MLE baseline has no regularization term, the change is significantly more abrupt.
In appendix B we also consider the discounted backward transfer and forgetting, using the formulae
$BWT^{\gamma}(Q_{1:T},[T])\triangleq\frac{1}{T-1}\sum_{t=1}^{T-1}\gamma^{T-1-t}\mathcal{L}(Q_{1:T},\mathcal{D}_{t}),$
$\begin{split}&F^{\gamma}(Q_{1:T},[T])\triangleq\\ &\frac{1}{T-1}\sum_{t=1}^{T-1}\gamma^{T-1-t}(\mathcal{L}(Q_{1:T},\mathcal{D}_{t})-\mathcal{L}(Q_{1:t}),\mathcal{D}_{t}).\end{split}$
These discounted forms allow us to re-weigh the forgetting based on recency, as achieving low forgetting on recent tasks may be more important in some settings than low overall forgetting. In general (for $\gamma=0.95$), the discounted backward transfer is lower, and the forgetting is higher. This is likely due to the shared model improving over time.
@@
Figure 2: Metrics over time for the permuted-MNIST (top row) and split-MNIST (bottom row) tasksets. (a) backward transfer and its upper bound (3.4) (b) forgetting and its upper bound (3.5) (c) average test error over time.
@@
In this section, we examine Algorithm 1 on more realistic problem domains, namely permuted-MNIST (Goodfellow et al., 2015), a domain-incremental continual learning problem (De Lange et al., 2021) where a random permutation of the pixels is applied on each image for each task, and split-MNIST (Zenke et al., 2017), a sequential set of binary classification tasks constructed from the MNIST dataset. We used convolutional neural networks for these visual domains, and $T=80$ tasks in total for both.
Figure 2() shows the gradual change in backward transfer. We can see that after an initial warm-up period, both settings have gradually improving backward transfer as well as tighter upper bounds (3.3). Figure 2() shows that for both settings, the backward transfer loss behaves similarly to the test error, also decreasing with additional tasks. This suggests that the shared representation has sufficient network capacity to learn a useful representation for all MNIST vision tasks.
Figure 2() shows the forgetting and its upper bound (3.4) over time. For the split-MNIST problem, we see similar behavior to that of the $10d$ Gaussian data, where the forgetting increases, decreases and then begins to trend back upwards toward zero. For the permuted-MNIST problem we only see a general upward trend. The difference coincides well with the change in test loss (forward transfer) over time: in the permuted-MNIST setting we see constant improvement whereas whereas for the split-MNIST setting we see a plateau in forward transfer until roughly $T=20$ that corresponds to an increase in forgetting, followed by a decrease in both forgetting and forward transfer as the shared representation improves.
The full tables of the results appear in Appendix B
@@ 6 Conclusions
In this work, we derived several upper bounds on the test forgetting (via backward transfer) for both general model classes and for the Gibbs posterior, based on the change of measure approach. These upper bounds are data-dependent, potentially offering tighter bounds if improved prior models for the task are available, or if tasks are structured such that their loss landscapes are similar. In particular, we focused on oracle bounds for Gibbs posteriors that offered tight bounds on backward transfer if task losses are highly positively correlated, thus making the knowledge accumulation process highly effective for all tasks.
@@
Based on our theoretical bounds, we constructed an algorithm for continual learning with empirically low forgetting and tight upper bounds. We examined this approach on several simple task constructions as well as slightly more complex vision tasks. In our experiments we noted a relation between forward and backward transfer, especially for mostly static settings such as the domain-incremental continual learning problem. As noted by several previous theoretical works and practical examinations (see Introduction), task order and similarity can greatly influence both forgetting and generalization. While this empirical demonstration is not the main focus of our paper, it suggests that weighting schemes based on notions of loss agreement merit further exploration for the domain incremental setting.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary, and section $2$ details the overall setting.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] Algorithm 1 has a short analysis comparing its time and sample complexity properties compared to standard variational inference without the novel upper bound.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Attached as a zip file as supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary, and section $2$ details the overall setting.
    2. (b)
      Complete proofs of all theoretical results. [Yes] Appendix A contains full proofs of all theoretical results.
    3. (c)
      Clear explanations of any assumptions. [Yes] General assumptions are clearly stated, specific assumptions are written for each theorem and corollary.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] Code (with data download) is attached as a zip file, readme file contains the flags used for reproduction.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] detailed in Appendix B.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] As listed, standard error over $5$ random seeds is used for all measures.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] Detailed in Appendix B.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] Datasets are cited.
    2. (b)
      The license information of the assets, if applicable. [Yes] Dataset licence (GNU GPLv3) is included.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes] Code is included in the supplemental material.
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Appendix A Appendix - proofs
Lemma A.1.

### v3

@@ 5 Empirical validation
In this section we demonstrate the utility of our bounds on both synthetic and real-world data sets. We study three classes of task environments, with very different types of shifts between tasks, and show, that for sufficiently many tasks, the bounds are tight. Since Corollary 3.2 applies uniformly over all posteriors, it can be optimized at each time step w.r.t. the current posterior. Specifically, we select
$\hat{Q}_{1:i}=\mathop{\mathrm{argmin}}_{Q_{1:i}}\left\{\hat{\mathcal{L}}(Q_{1:i},S_{i})+\frac{1}{\lambda_{i}}D_{\mathrm{KL}}(Q_{1:i}||Q_{1:i-1})\right\}.$ (8)
This optimization process yields a theoretically motivated stochastic optimization algorithm, detailed in Algorithm 1.
@@
4:for each task $t$ from 11 to $T$ do
@@
This algorithm can be approximated in practice via variational inference (VI) using multivariate Gaussian prior and posterior distributions on model parameters. The main benefit of doing so is to simplify the estimation of the KL-divergence. Other distributions are possible, but Gaussian distributions are common in VI (Hoffman et al., 2013). The posterior can be used to estimate the forgetting and its upper bound by using a pre-defined test set as a proxy for the data distribution $\mathcal{D}_{i}$. Calculating the upper bound (3.2) would add an additional calculation of $O(t\cdot(m+d))$ to each task, where $m=|S_{t}|$ and $d$ is the number of model parameters. In total, calculating the bound at each task will add $O(T^{2}\cdot(m+d))$. Calculating the bound does not require additional memory or samples (samples from previous test tasks are also required for calculating test forgetting).
We also examine our bounds for a deterministic continual learning algorithm, namely Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017). Since EWC outputs deterministic parameters for each task, a posterior distribution is constructed by adding Multivariate Gaussian noise, i.e. $\hat{Q}_{1:i}=\mathcal{N}(w_{i},\sigma^{2}I_{d})$, where $w_{i}\in\mathcal{R}^{d}$ is the weight vector given as output after task $\mathcal{D}_{i}$. By doing so, the KL-divergence can be easily calculated as
$KL(\hat{Q}_{1:i}||\hat{Q}_{1:i-1})=\frac{1}{2\sigma^{2}}||w_{i}-w_{i-1}||^{2}_{2}.$
@@
In order to measure the bound tightness for VI and EWC, we must define the relevant metrics we wish to measure. These are the backwards transfer $BWT(\hat{Q}_{1:T},[T])$ (measured via a separate test set per task) and the forgetting $F(\hat{Q}_{1:T},[T])$, as well as the average forward transfer $\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}(\hat{Q}_{1:t},\mathcal{D}_{t})$.
In the following experiments, we use fully connected neural networks with a separate linear head per task to facilitate measuring the forgetting without re-training linear heads. We use Adam (Kingma & Ba, 2015) for optimization. The full list of hyper-parameters is listed in Appendix C.
@@
We begin by examining several setups of binary classification tasks in $\mathbb{R}^{10}$. All tasks draw samples from a $10$-dimensional Gaussian distribution $p(x)=\mathcal{N}(x;0,I_{10})$ and $y=\mathrm{sgn}(a^{\top}x)$. To simplify, only the first two features were used to determine $y$, meaning that there is a $2d$ linear separator embedded in $\mathbb{R}^{10}$. We consider the following settings: (1) Similar tasks: linear separators for all tasks are within $10^{\circ}$ of some reference angle. (2) Gradual shift: tasks change angle gradually in a set direction, with each consecutive task being within $10^{\circ}$ of the previous one. (3) Orthogonal shift: Tasks originate from two distinct, orthogonal directions, with the first half of tasks being from the first angle and the latter half corresponding to the second angle.
We used a total of $T=100$ tasks in this domain in order to observe general trends. We note that these problems differ significantly in task order and overall behavior, and are aimed at providing a diverse set of challenges. In addition, we should not expect a single algorithm to perform optimally across all settings, as known empirical methods based on regularization would likely perform poorly on orthogonal tasks whereas replay-based and architecture-based methods would have issues with gradual shift. We expect that in real-world settings, prior knowledge can be used to suggest specific algorithms. In particular, the settings of Gradual shift and Orthogonal shifts are constructed such that forgetting may be desirable. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially given changing environments. In both the gradual and orthogonal shift settings, not all previous knowledge is beneficial.
@@
Method BWT BWT bound Forgetting Forgetting bound (3.2) Test loss
@@
EWC $2.0\pm 0.1$ $3.5\pm 0.1$ $0.4\pm 0.0$ $1.8\pm 0.1$ $1.8\pm 0.4$
VI $\textbf{0.8}\pm 0.2$ $\textbf{2.2}\pm 0.2$ -$\textbf{-1.9}\pm 0.1$ -$\textbf{-0.5}\pm 0.1$ $\textbf{1.0}\pm 0.5$
@@
EWC $1.9\pm 0.1$ $3.3\pm 0.1$ $0.3\pm 0.1$ $1.7\pm 0.1$ $2.1\pm 0.2$
VI $\textbf{0.2}\pm 0.0$ $\textbf{1.5}\pm 0.0$ -$\textbf{-2.2}\pm 0.0$ -$\textbf{-0.8}\pm 0.0$ $\textbf{0.2}\pm 0.1$
@@
EWC $4.1\pm 0.8$ $5.5\pm 0.9$ $0.8\pm 0.2$ $2.2\pm 0.2$ $4.6\pm 0.9$
VI $\textbf{3.3}\pm 0.8$ $\textbf{4.7}\pm 0.8$ -$\textbf{-0.0}\pm 0.7$ $\textbf{1.4}\pm 0.7$ $\textbf{0.5}\pm 0.1$
@@
(b) Orthogonal shift at $t=50$
Figure 2: Forgetting and upper bound (3.2) over time ($t=4$ to $t=96$) for $10d$ data.
Table 2 describes the forward and backward transfer as well as forgetting for the $10d$ tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign. Best result for each setting is in bold. We can see that for most settings, VI based on Algorithm 1 tends to provide good forward transfer, as well as backward transfer and forgetting. For the orthogonal shift setting, we can see that both backward transfer and forgetting degrade. Figure 2 compares the forgetting for the gradual shift setting compared to the more abrupt orthogonal shift at $T=50$. We can clearly see that the backward transfer increases abruptly at the point of task shift, though the upper bound remains roughly the same in terms of tightness as the task discrepancy term is negative and corrects for the higher KL-divergence at the task shift.
The forgetting for VI is negative and seems to follow a trend of increasing gradually towards zero forgetting. For the orthogonal shift setting, the forgetting becomes positive and decreases towards zero. For EWC, the forgetting is positive and decreases gradually towards zero. We observe that as the shared model improves, forgetting decreases, likely due to the learned shared representation improving overall across several tasks, and we see the forgetting trend towards zero. An interesting observation is that the backward transfer (and forgetting) for the orthogonal shift setting is much higher. In this setting, forgetting may be a desired phenomenon, as the data distribution shifts in the middle of the learning process. We see that indeed this shift leads to higher forgetting.
In appendix C we also consider the discounted backward transfer and forgetting, using the formulae
$\begin{split}&BWT^{\gamma}(Q_{1:T},[T])\triangleq\sum_{t=1}^{T-1}\gamma^{T-1-t}\mathcal{L}(Q_{1:T},\mathcal{D}_{t}),\quad F^{\gamma}(Q_{1:T},[T])\triangleq\sum_{t=1}^{T-1}\gamma^{T-1-t}(\mathcal{L}(Q_{1:T},\mathcal{D}_{t})-\mathcal{L}(Q_{1:t}),\mathcal{D}_{t}).\end{split}$
These discounted forms allow us to re-weigh the forgetting based on recency, as achieving low forgetting on recent tasks may be more important in some settings than low overall forgetting. For $\gamma=0.95$, the discounted backward transfer for VI is lower, and the forgetting is higher. This is likely due to the shared model improving over time. For EWC, the backward transfer increases until a low steady state, and forgetting decreases gradually.
@@
Figure 3: Metrics over time for permuted-MNIST (top row), split-MNIST (middle row) and split-CIFAR10 (bottom row). (a) backward transfer & upper bound (b) forgetting & upper bound (3.2) (c) average (forward) test error over time.
@@
Next we examine our bounds for our approximation of Algorithm 1 and EWC on more realistic problem domains, namely permuted-MNIST (Goodfellow et al., 2015), a domain-incremental continual learning problem (De Lange et al., 2021) where a random permutation of the pixels is applied on each image for each task, split-MNIST (Zenke et al., 2017), a sequential set of binary classification tasks constructed from the MNIST dataset, and split-CIFAR10 (Zenke et al., 2017), a sequential set of binary classification tasks constructed from the CIFAR-10 dataset. We used convolutional neural networks for these visual domains, with $T=120$ tasks in total for MNIST domains and $T=100$ for split-CIFAR10. For these more challenging tasks, our practical approximation of Algorithm 1, that works well for synthetic data, does not perform as well as EWC in terms of forward and backward transfer.
Figure 3() shows the gradual change in backward transfer. We can see that after an initial warm-up period, all settings have gradually improving backward transfer as well as tight upper bounds (16), especially for EWC. Figure 3() shows that for both MNIST settings, the test error for VI decreases with additional tasks. This suggests that the shared representation has sufficient network capacity to learn a useful representation for all MNIST vision tasks. For EWC the loss also decreases, but rapidly reaches near zero. For the split-CIFAR10 setting, we see that VI fails to learn useful hypotheses, and EWC performs reasonably well.
Figure 3() shows the forgetting and its upper bound (3.2) over time. For both MNIST problems, we see similar behavior to that of the $10d$ Gaussian data, where the forgetting trends upwards toward zero. For split-CIFAR10 we see the forgetting for EWC remain at a mostly constant low level as the backward transfer and test losses improve slowly. We see that the bounds for split-CIFAR10 are reasonably tight. The poor performance of VI compared to EWC on the split-CIFAR10 can be attributed to a combination of low computational budget and modeling choices. VI methods often require additional computation epochs compared to deterministic neural networks, and computational parity with EWC was prioritized. It may be the case that the approximation via an isotropic Gaussian prior served to limit the search space for optimization. The full tables of the results appear in Appendix C.
6 Limitations
Our upper bounds assume that tasks are sampled independently of the learning process. This assumption is necessary for our proofs and while reasonable for supervised learning (outside of curriculum learning), would not hold under adversarial task constructions. We also assume that the loss function $\ell(h,z)$ is task-independent. This assumption can likely be relaxed via separate loss functions per task without invalidating the proof. Our bounds assume that the the continual learner is exemplar-free (data from previous tasks is not directly available). It may be the case that additional data availability may be used to construct tighter bounds. Additional assumptions are listed throughout the text, with many of our results assuming that the loss is bounded (Assumption 1). Most of these results can be extended to low-tailed losses, such as sub-Gaussian losses, but would not hold for heavy-tailed losses.
Our empirical results are mainly focused on measuring the tightness of our upper bounds. As such, we focused on relatively small neural networks and reasonably simple supervised problems in order to validate our bounds in practice. Since the KL-divergence is part of the upper bound (3.2), it may be the case that this bound will become looser for larger models, as the KL-divergence tends to increase with the number of parameters. Compared to standard PAC-Bayes bounds this issue is somewhat alleviated due to only requiring the KL-divergence for the final model, meaning that this term can be mitigated somewhat by ensuring that model changes lessen over time. Outside of our synthetic dataset, all of our experiments used vision datasets, and we only examined our bounds on our simple stochastic algorithm and EWC, and therefore our experiments should be considered preliminary in scope.
@@ 7 Conclusions
In this work, we derived several upper bounds on the test forgetting (via backward transfer) for both general model classes and for the Gibbs posterior, based on the change of measure approach. These upper bounds are data-dependent, offering tighter bounds if improved prior models for the task are available, or if the loss landscapes for tasks are similar. In particular, we focused on oracle bounds for Gibbs posteriors that offered tight bounds on backward transfer if task losses are highly positively correlated, thus making the knowledge accumulation process highly effective for all tasks.
@@
Based on our theoretical bounds, we constructed an algorithm for continual learning with empirically low forgetting and tight upper bounds. We examined our bounds for both this algorithm and EWC on several simple task constructions as well as slightly more complex vision tasks. In our experiments, we noted a relation between forward and backward transfer, especially for mostly static settings such as the domain-incremental continual learning problem. As noted by several previous theoretical works and practical examinations (see Introduction), task order and similarity can greatly influence both forgetting and generalization. While this empirical demonstration is not the main focus of our paper, it suggests that weighting schemes based on notions of loss agreement merit further exploration for the domain incremental setting.
Acknowledgements
This work was supported by the Israel Science Foundation grant number $1693/22$.
Appendix A Appendix - notation glossary
$\mathcal{X}$ Feature set
$x\in\mathcal{X}$ Data sample
$\mathcal{Y}$ Label set
$y\in\mathcal{Y}$ Data label
$T$ Number of tasks
$\mathcal{D}$ Task distribution over pairs $(x,y)$
m Size of training set
$S\sim\mathcal{D}^{m}$ Training set of size $m$
$\mathcal{H}$ Hypothesis set
$h\in\mathcal{H},h:\mathcal{X}\rightarrow\mathcal{Y}$ Hypothesis
$\ell:\mathcal{H}\times(\mathcal{X},\mathcal{Y})\rightarrow\mathbb{R}$ Loss function
$\mathcal{L}:\mathcal{H}\times\mathcal{D}\rightarrow\mathbb{R}$ Expected loss
$\hat{\mathcal{L}}:\mathcal{H}\times(\mathcal{X},\mathcal{Y})^{m}\rightarrow\mathbb{R}$ Empirical loss
$P$ Data-free prior distribution over hypotheses
$Q$ Data-dependent posterior distribution over hypotheses
$Q_{1:i}$ Data-dependent posterior distribution depending on $S_{1},\ldots,S_{i}$
$\lambda>0$ KL-divergence weight (Gibbs posterior temperature)
$K$ Maximal loss value
$\delta\in(0,1)$ A probability
$d$ Dimension of hypothesis space under Assumption 2
$\sigma^{2}$ Variance (Gaussian distribution)
Appendix B Appendix - proofs
Lemma B.1.